1. a) Show or verify that 3 is a primitive root of the prime p=7.

b) Show that 2 is not a primitive root of 7.

c) Show the steps of the Diffie-Hellman key agreement protocol between Alice and Bob, assuming they use the prime 7 and its primitive root 3, and Alice’s secret integer is k1=5 and Bob’s secret integer is k2=4.


3 Consider the function f(n) = 2n2-5n+1.

a) What is the smallest value of n for which f(n)0 ?

b) Use mathematical induction to prove that f(n)0 for all n the value in part a). 


4 Write the pseudocode for a recursive algorithm to compute b3k, where b is a real number and k is a positive integer.  Use the fact that b3k+1=(b3k)3.







1
Expert's answer
2021-12-02T10:16:11-0500

1.

a)

If p is prime, then b is a primitive root for p if the powers of b include all of the residue classes mod p



Since we achieved all values from 1 to 6 in our residue results, then 3 is a primitive root of 7


b)

A primitive root modulo 7 would have order 6, but "2^3=8\u22611(mod7)" ,

so 2 is not a primitive root modulo 7.


3.

a)

"2x^2-5x+1=0"

"x=\frac{5\pm\sqrt{25-8}}{4}=\frac{5\pm\sqrt{17}}{4}"

"x_1=0.22,x_2=5.15"

"f(x)=2(x-0.22)(x-5.15)"

"f(0)>0" for "x<0.22" and "x>5.15"

smallest value of n > 0 for which f(n) > 0 :

"n=6"


b)

for n=6:

"f(6)=43>0"

for n=k let:

"2k^2-5k+1>0"

then for n=k+1:

"2(k+1)^2-5(k+1)+1=2k^2-5k+1+4k-3>0"

since "2k^2-5k+1>0" , and "4k-3>0" for "k>6"


4.

input:

b

int k

f(0)=b


output:

f(k+1)=3*f(k)


1.

c)

Alice and Bob agree to use the prime p = 7 and the primitive root g = 3. Alice chooses the secret key k1 = 5 and computes

"A=5\equiv 3^5(mod\ 7)"

Bob chooses the secret key k2 = 4 and computes

"B=4\equiv 3^4(mod\ 7)"

Alice sends Bob the number 5 and Bob sends Alice the number 4. Both of these transmissions are done over an insecure channel, so both A = 5 and B = 4 should be considered public knowledge. The numbers k1 = 5 and k2 = 4 are not transmitted and remain secret. Then Alice and Bob are both able to compute the number

"3^{k_1\cdot k_2}\equiv A^{k_2}\equiv B^{k_1}(mod\ 7)"

"2\equiv3^{5\cdot 4}\equiv 5^{4}\equiv 4^{5}(mod\ 7)"

so 2 is their shared secret.


Suppose that Eve sees this entire exchange. She can reconstitute Alice’s and Bob’s shared secret if she can solve either of the congruences 

"3^{k_1}\equiv 5(mod\ 7)" or "3^{k_2}\equiv 4(mod\ 7)"

since then she will know one of their secret exponents.